OpenStudy (anonymous):

How do I solve? y<= –x + 7 y < 3x + 11

OpenStudy (anonymous):

More correctly speaking, how do i set it up to solve?

OpenStudy (anonymous):

Hm, heard about linear programming ?

OpenStudy (anonymous):

This is completely new to me

OpenStudy (anonymous):

The graphical solution is trivial and probably the intended one, just bring them to slope intercept form and plot the lines .. shade the respective region and try to identify the intersection.

OpenStudy (anonymous):

i can bring the second equation to y=mx+b how how do i do it for the first one?

OpenStudy (anonymous):

y=mx+b is not the slope intercept form x/a+y/b=1 is, also basic rules of inequality applies when you are modifying.

OpenStudy (anonymous):

I am so lost... my text book doesnt like to clarify things... it told me that slope intercept was y=mx+b...

OpenStudy (anonymous):

Oops my fault, x/a+y/b=1 is only the intercept form.

OpenStudy (anonymous):

ok... what would be a in this problem. i am really sorry to bother you, my textbook is all i have for reference and it told us to work it out and if we couldnt figure it out ask the teacher.. the teacher doesnt know...

OpenStudy (anonymous):

a is the x intercept and b is the y intercept.

OpenStudy (anonymous):

what would be the first step? 3x/a+y/b=1?

OpenStudy (anonymous):

\( y < 3x + 11\Rightarrow \frac y{11} + \frac x {\frac {11} 3} \lt 1 \)

OpenStudy (anonymous):

Try to do the other one yourself.

OpenStudy (anonymous):

y/7+x/11/3<=1?

OpenStudy (anonymous):

Yes. now plot them :)

OpenStudy (anonymous):

sorry x/7/3

OpenStudy (anonymous):

\( y \le –x + 7 \Rightarrow \frac y7 + \frac x7 \le 1 \)

OpenStudy (anonymous):

wow... i guess this shows that i am getting tired... thanks... so the two equations are: \[y/7 +x/7\le1\] \[y/11+x/11/3<1\] So I find the LCM and solve?

OpenStudy (anonymous):

Dude, don't you know anything about graphing inequalities ?

OpenStudy (anonymous):

No. This is the first I am hearing... Thanks to my school they should change their motto to we wont teach you so dont bother... but I want to learn...

OpenStudy (anonymous):

How old is you ?

OpenStudy (anonymous):

This should help : http://www.purplemath.com/modules/ineqgrph.htm

OpenStudy (anonymous):

I am 14.

OpenStudy (anonymous):

Great, so much time to learn mathematics! :)

OpenStudy (anonymous):

Yes, but no one at my school wants to teach.

OpenStudy (anonymous):

I know, you will find this interesting http://www.maa.org/devlin/LockhartsLament.pdf

OpenStudy (anonymous):

what is it?

OpenStudy (anonymous):

Some English words joined together to form meaningful sentences .. ;)

OpenStudy (anonymous):

It's will explain you why teacher don't teach you and why some people hate maths.

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

where do i know to put the slope line?